OpenStudy (anonymous):

if f(x)={(3,5),(2,4),(1,7)} and g(x)=radical(x-3) what do i do to determine (f+g)(1)

OpenStudy (anonymous):

(f+g)(1) is the same as f(1)+g(1). f(1) is given from the set for f(x). The ordered pair (1,7) means f(1)=7. g(1) requires calculation from g(x). \[g(1)=\sqrt{1-3}=\sqrt{-2}=i \sqrt2\] That answer isn't a real number, so there's two possible answers. If you're only concerned with real numbers, (f+g)(1) is undefined. If you're looking at complex numbers, \[(f+g)(1)=f(1)+g(1)=7+i \sqrt2\]